# Emergent Weyl-like points in periodically modulated systems and Floquet engineering

### v3

@@ II.2 Weyl semimetals
where ${\bf k}\!=\!(k_{x},k_{y},k_{z})$. This Hamiltonian describes a Weyl semimetal phase and can be viewed as a stack of 2D layers with orbital-dependent interlayer hopping $t_{z}$, as illustrated in Fig. 1(a2). The matrix representation of Eq. (10) in real space with open boundary conditions (OBCs) along the $z$-direction and periodic boundary conditions (PBCs) along the $x$- and $y$-directions is provided in Subsection SIII A of the Supplemental MaterialSup . Similarly, the matrix form with OBCs in the $y$-direction and PBCs in the $x$- and $z$-directions is given in Subsection SIII B in the Supplemental Material Sup .
We can rewrite Eq. (10) as ${\cal H}_{\rm W}({\bf k})\!=\!\mathbf{h}({\bf k})\cdot{\boldsymbol{\sigma}}$, where the vector $\mathbf{h}({\bf k})\!=\!(h_{x},h_{y},h_{z})$ has the following components: $h_{x}\!=\!t_{x}\sin(k_{x}a)$, $h_{y}\!=\!t_{y}\sin(k_{y}a)$, $h_{z}\!=\!M_{z}(k_{z})\!-\!m_{0}\left[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)\right]$, with $M_{z}(k_{z})\!=\!-m_{z}\!+\!t_{z}\cos(k_{z}a)$. The eigenenergies of the Weyl semimetal model then take the form
@@
The Chern number of this model (10) at fixed $k_{z}$ can be calculated using the expression Thouless et al. (1982); Fukui et al. (2005); Xiong et al. (2016); Liu et al. (2019); Lu et al. (2010); Shan et al. (2010); Qin et al. (2022a)
$C(k_{z})\!=\!\frac{1}{4\pi}\int_{\text{B.Z.}}\!\!d^{2}{\bf k}_{||}\frac{\mathbf{h}({\bf k})}{|\mathbf{h}({\bf k})|^{3}}\!\cdot\![\partial_{k_{x}}\mathbf{h}({\bf k})\!\times\!\partial_{k_{y}}\mathbf{h}({\bf k})],$
where the integral is over the 2D Brillouin zone. A detailed analytical expression for the Chern number (II.2) is provided in Section SIV in the Supplemental Material Sup .
@@
$E_{\pm}\!=\!\pm\left[-m_{z}\!+\!t_{z}\cos(k_{z}a)\right].$ (13)
so the corresponding bulk energy gap at $k_{x}\!=\!k_{y}\!=\!0$ is given by $\Delta_{\rm W}\!=\!E_{{\rm W},+}\!-\!E_{{\rm W},-}\!=\!2|M_{z}(k_{z})|$ as shown in Fig. 1(d2). In this case, the system hosts a pair of gapless Weyl nodes at $(0,0,k_{c\pm})$ with the positions $k_{c\pm}\!=\!\pm\frac{1}{a}\arccos(m_{z}/t_{z})$, under the condition $|m_{z}/t_{z}|\leqslant 1$ as shown in Figs. 1(b2) and 1(c2). These Weyl points $k_{c\pm}$ are are determined by solving the condition $M_{z}(k_{z})\!=\!-m_{z}\!+\!t_{z}\cos(k_{z}a)\!=\!0$. For $k_{z}$ in the range $k_{c-}<k_{z}<k_{c+}$, the system behaves as a Chern insulator with Chern number $|C|\!=\!1$. Outside this range, for $k_{z}<k_{c-}$ or $k_{z}>k_{c+}$, the system becomes a normal insulator with $C\!=\!0$, as depicted in Fig. 1(c2). The corresponding Fermi arc is shown in Fig. 1(e2).
Finally, the total 3D Hall conductivity for the Weyl semimetal model (10) is given by Burkov and Balents (2011)
@@
where $\sigma_{H}^{2D}(k_{z})\!=\!(e^{2}/h)C(k_{z})$ is the $k_{z}$-resolved 2D Hall conductance and $C(k_{z})$ is defined in Eq. (II.2). Using the Weyl node positions $k_{c\pm}\!=\!\pm\frac{1}{a}\arccos(m_{z}/t_{z})$, we obtain
@@
Figure 2: (a) Weyl-like nodes at $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$, marking topological transitions. (b) The two Weyl-like nodes are separated by the bulk gap between $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$. Each half of the system contains a single Weyl-like node. This differs from conventional Weyl points, which typically appear in pairs in momentum space $(k_{x},k_{y},k_{z})$.
@@ II.3 Comparison between the two models
As shown in Fig. 1, the two models – the decoupled Weyl-like model (left column) and the Weyl semimetal model (right column) – can be viewed as duals of each other. The Weyl-like model in Eq. (1) provides a generic framework for describing Weyl semimetals, including key features such as band touching, Chern number, and Fermi arc, all of which are captured by the Weyl semimetal model in Eq. (10).
@@
Figure 2 illustrates that the two Weyl-like nodes are spatially separated by a bulk gap between $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$, with each half of the system hosting a single Weyl-like node. This contrasts with conventional Weyl points, which, due to the no-go theorem Nielsen and Ninomiya (1981a, b), typically appear in pairs in momentum space with opposite chiralities. In the next section, we turn our attention to the effects of interlayer coupling on the Weyl-like model Hamiltonian (1).
@@ III Interlayer coupling effects
We now incorporate the effects of interlayer coupling into our analysis. The energy spectrum is still obtained by calculating the local density of states Chen et al. (2021); Wan et al. (2011); Yang et al. (2011), while the quantum transport properties are captured by evaluating the layer Hall conductance Fukui et al. (2005); Sun et al. (2020); Otrokov et al. (2019); Gao et al. (2021); Chen et al. (2024a); Mong et al. (2010); Essin et al. (2009); Wang et al. (2015); Varnava and Vanderbilt (2018); Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019) in the presence of interlayer coupling. Additionally, we examine how interlayer coupling influences the total 3D Hall conductivity.
@@
where $c_{j_{z}}^{{\dagger}}$ ($c_{j_{z}}$) is the electron creation (annihilation) operator on the $j_{z}$-th layer and $\eta_{z}\sigma_{z}$ represents the nearest-neighbor interlayer hopping matrix. Combining Eq. (1) with the interlayer hopping term, the full Hamiltonian incorporating interlayer coupling becomes
@@
The matrix representation of the Hamiltonian (17) under OBCs along $z$-direction is provided in Subsection SV A in the Supplemental Material Sup .
@@ III.1 Local density of states
Figure 3: Local density of states (LDOS) [Eq. (18)] for different interlayer hopping strengths $\eta_{z}$. The LDOS is indicated by the red color, while the two green vertical dashed lines mark the positions of the Weyl-like points. (a) $\eta_{z}/t_{0}\!=\!0$; (b) $\eta_{z}/t_{0}\!=\!0.05$; (c) $\eta_{z}/t_{0}\!=\!0.1$; (d) $\eta_{z}/t_{0}\!=\!0.2$. The other parameters are $k_{x}\!=\!k_{y}\!=\!0$, $k_{B}T=0.01$ eV, $V_{z}\!=\!1$ eV, $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$ with $a\!=\!1$ nm.
@@
$g(E,E_{n},T)$ is a Lorentzian broadening function defined as
@@
Here, $E$ is the probe energy, $k_{B}$ is the Boltzmann constant, and $T$ is the temperatrue.
@@
For weak interlayer coupling, $\eta_{z}/t_{0}\!=\!0.05$, Fig. 3(b) reveals that the local density of states still exhibits sharp crossings at the same Weyl-like points, similar to the decoupled case in Fig. 3(a). For moderate coupling, $\eta_{z}/t_{0}\!=\!0.1$, Fig. 3(c) shows that the local density of states continues to display distinct crossings at the Weyl-like points, indicating that the features observed in the decoupled case persist. These results suggest that the Weyl-like points are robust against weak to moderate interlayer coupling. However, as the interlayer hopping strength $\eta_{z}$ increases further, Fig. 3(d) shows that the local density of states becomes increasingly dispersive compared to the decoupled scenario. This behavior indicates that strong interlayer coupling disrupts the localized character of the Weyl-like points, ultimately leading to their destruction.
@@ III.2 Layer Hall conductance
Figure 4: Layer Hall conductances for different interlayer hopping strengths $\eta_{z}$. The two green vertical dashed lines indicate the positions of the Weyl-like points. (a) $\eta_{z}/t_{0}\!=\!0$; (b) $\eta_{z}/t_{0}\!=\!0.05$; (c) $\eta_{z}/t_{0}\!=\!0.1$; (d) $\eta_{z}/t_{0}\!=\!0.2$; (e) $\eta_{z}/t_{0}\!=\!0.4$; (f) $\eta_{z}/t_{0}\!=\!0.6$. All other parameters are the same as those used in Fig. 3.
@@
However, as the interlayer hopping strength $\eta_{z}$ increases further [Figs. 4(d)–4(f)], the layer Hall conductance becomes increasingly smooth and continuous in the vicinity of the Weyl-like points, deviating from the sharp features observed in the decoupled case. Physically, this behavior indicates that strong interlayer coupling destroys the Weyl-like points and significantly alters the bulk topology of the system.

### v4

@@ II.2 Weyl semimetals
where ${\bf k}\!=\!(k_{x},k_{y},k_{z})$. This Hamiltonian describes a Weyl semimetal phase and can be viewed as a stack of 2D layers with orbital-dependent interlayer hopping $t_{z}$, as illustrated in Fig. 1(a2). The matrix representation of Eq. (10) in real space with OBCs along the $z$ direction and PBCs along the $x$ and $y$ directions is provided in Subsection SIII A in the Supplemental Material Sup . Similarly, the matrix form with OBCs in the $y$ direction and PBCs in the $x$ and $z$ directions is given in Subsection SIII B in the Supplemental Material Sup .
We can rewrite Eq. (10) as ${\cal H}_{\rm W}({\bf k})\!=\!\mathbf{h}({\bf k})\cdot{\boldsymbol{\sigma}}$, where the vector $\mathbf{h}({\bf k})\!=\!(h_{x},h_{y},h_{z})$ has the following components: $h_{x}\!=\!t_{x}\sin(k_{x}a)$, $h_{y}\!=\!t_{y}\sin(k_{y}a)$, and $h_{z}\!=\!M_{z}(k_{z})\!-\!m_{0}\left[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)\right]$, with $M_{z}(k_{z})\!=\!-m_{z}\!+\!t_{z}\cos(k_{z}a)$. The eigenenergies of the Weyl semimetal model then take the form
@@
The Chern number of this model in Eq. (10) at fixed $k_{z}$ can be calculated using the expression Thouless et al. (1982); Fukui et al. (2005); Xiong et al. (2016); Liu et al. (2019); Lu et al. (2010); Shan et al. (2010); Qin et al. (2022a)
$C(k_{z})\!=\!\frac{1}{4\pi}\int_{\text{BZ}}\!\!d^{2}{\bf k}_{||}\frac{\mathbf{h}({\bf k})}{|\mathbf{h}({\bf k})|^{3}}\!\cdot\![\partial_{k_{x}}\mathbf{h}({\bf k})\!\times\!\partial_{k_{y}}\mathbf{h}({\bf k})],$
where the integral is over the 2D Brillouin zone (BZ). A detailed analytical expression for the Chern number in Eq. (II.2) is provided in Section SIV in the Supplemental Material Sup .
@@
$E_{\pm}\!=\!\pm\left[-m_{z}\!+\!t_{z}\cos(k_{z}a)\right],$ (13)
so the corresponding bulk energy gap at $k_{x}\!=\!k_{y}\!=\!0$ is given by $\Delta_{\rm W}\!=\!E_{{\rm W},+}\!-\!E_{{\rm W},-}\!=\!2|M_{z}(k_{z})|$, as shown in Fig. 1(d2). In this case, the system hosts a pair of gapless Weyl nodes at $(0,0,k_{c\pm})$ with the positions $k_{c\pm}\!=\!\pm(1/a)\arccos(m_{z}/t_{z})$, under the condition $|m_{z}/t_{z}|\leqslant 1$, as shown in Figs. 1(b2) and 1(c2). These Weyl points $k_{c\pm}$ are determined by solving the condition $M_{z}(k_{z})\!=\!-m_{z}\!+\!t_{z}\cos(k_{z}a)\!=\!0$. For $k_{z}$ in the range $k_{c-}\!<\!k_{z}\!<\!k_{c+}$, the system behaves as a Chern insulator with Chern number |C|=1|C|\!=\!1. Outside this range, for $k_{z}\!<\!k_{c-}$ or $k_{z}\!>\!k_{c+}$, the system becomes a normal insulator with C=0C\!=\!0, as depicted in Fig. 1(c2). The corresponding Fermi arc is shown in Fig. 1(e2).
Finally, the total 3D Hall conductivity for the Weyl semimetal model in Eq. (10) is given by Burkov and Balents (2011)
@@
where $\sigma_{H}^{2D}(k_{z})\!=\!(e^{2}/h)C(k_{z})$ is the $k_{z}$-resolved 2D Hall conductance and $C(k_{z})$ is defined in Eq. (II.2). Using the Weyl node positions $k_{c\pm}\!=\!\pm(1/a)\arccos(m_{z}/t_{z})$, we obtain
@@
Figure 2: (a) Weyl-like nodes at $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$, marking topological transitions. (b) The two Weyl-like nodes are spatially separated by the bulk gap between $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$. Each half of the system contains a single Weyl-like node. This differs from conventional Weyl points, which typically appear in pairs in momentum space $(k_{x},k_{y},k_{z})$.
@@ II.3 Comparison between the two models
As shown in Fig. 1, the two models, the decoupled Weyl-like model (left column) and the Weyl semimetal model (right column), can be viewed as duals of each other. The Weyl-like model in Eq. (1) provides a generic framework for describing Weyl semimetals, including key features such as band touching, Chern number, and Fermi arc, all of which are captured by the Weyl semimetal model in Eq. (10).
@@
The Weyl-like nodes do not obey the no-go theorem that constrains ordinary Weyl points. According to the no-go theorem Nielsen and Ninomiya (1981a, b), Weyl points always appear in pairs. However, Fig. 2 illustrates that the two Weyl-like nodes are spatially separated by a bulk gap between $z\!=\!z_{c-}$ and $z\!=\!z_{c+}$, with each half of the system hosting a single Weyl-like node. This contrasts with conventional Weyl points, which, due to the no-go theorem Nielsen and Ninomiya (1981a, b), typically appear in pairs in momentum space with opposite chiralities. In the next section, we turn our attention to the effects of interlayer coupling on the Weyl-like model Hamiltonian in Eq. (1).
By comparing Eqs. (9) and (15), we see that the 3D Hall conductivity in the Weyl-like model explicitly depends on the system size $L_{z}$, while that of the Weyl semimetal model is independent of system size. As such, the finite-size dependence of the 3D Hall conductivity cannot be captured by the Weyl semimetal model. This unique feature provides a clear signature of mixed-space topology in the Weyl-like model.
II.4 Local Berry curvature
In this subsection, we calculate the local Berry curvature of the Weyl-like model in Eq. (1).
To present the vector plot of the Berry curvature in the mixed space $(k_{x},k_{y},z)$, we first define the Berry curvature vector as Xiao et al. (2010); Li et al. (2016); Lu and Shen (2017)
${\bf F}^{mn}({\bf k}_{||},z)\!=\!\nabla\!\times\!{\bf A}^{mn}({\bf k}_{||},z),$ (16)
where ${\bf A}^{mn}({\bf k}_{||},z)\!=\!i\langle\psi_{m}|\nabla\psi_{n}\rangle\!=\!({\cal A}_{k_{x}}^{mn},{\cal A}_{k_{y}}^{mn},{\cal A}_{z}^{mn})$ is the Berry connection vector in the mixed space $(k_{x},k_{y},z)$.
Figure 3: Vector plot of the Berry curvature ${\bf F}({\bf k}_{||},z)$ [Eq. (16)] (summed over all occupied states $m,n$) in the mixed space $(k_{x},z)$ with $k_{y}\!=\!0$ fixed for simplicity. A pair of Berry monopoles Li et al. (2016); Lu and Shen (2017) is located at the two Weyl-like nodes. Arrows indicate the flux of the Berry curvature, flowing from one monopole $(k_{x}\!=\!0,z_{c-}\!=\!10\frac{1}{3}a)$, blue point] to the other $(k_{x}\!=\!0,z_{c+}\!=\!20\frac{2}{3}a)$, black point], revealing the nontrivial topological nature of the Weyl-like model in Eq. (1). The parameters used are $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$.
Figure 3 shows the vector plot of the Berry curvature ${\bf F}({\bf k}_{||},z)$ in the mixed space $(k_{x},z)$, with $k_{y}\!=\!0$ fixed for simplicity. A pair of Berry monopoles Li et al. (2016); Lu and Shen (2017) is located at the two Weyl-like nodes. The arrows indicate the flux of the Berry curvature, which flows from one monopole $(k_{x}\!=\!0,z_{c-}\!=\!10\frac{1}{3}a)$, blue point] to the other $(k_{x}\!=\!0,z_{c+}\!=\!20\frac{2}{3}a)$, black point], reflecting the nontrivial topological nature of the Weyl-like model in Eq. (1).
The chirality of a Weyl-like node can be obtained by integrating the Berry curvature over a Fermi surface enclosing the node Li et al. (2016); Lu and Shen (2017). This procedure yields opposite topological charges for the two Weyl-like points.
Figure 4: Momentum distribution of the layer Berry curvature ${\cal F}_{xy}({\bf k}_{||},z)$ [Eq. (4)] (summed over all occupied states $m,n$) in the $(k_{x},k_{y})$ plane for different layers: $j_{z}\!=\!z/a\!=\!10$, $11$, 2020, and 2121 (from left to right). (a) z=10az\!=\!10a, (b) z=11az\!=\!11a, (c) z=20az\!=\!20a, and (d) z=21az\!=\!21a. Here, $a$ is the lattice constant, and ${\cal F}_{xy}$ denotes the layer Berry curvature. The two critical planes are located at $z_{c-}\!=\!10\frac{1}{3}a$ and $z_{c+}\!=\!20\frac{2}{3}a$, given by $z_{c\pm}\!=\!\frac{\lambda_{z}}{2\pi}\left[\pi\pm\arccos\left(\frac{m_{z}}{V_{z}}\right)\right]$ [see Eq. (7)]. The parameters used are $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$.
Furthermore, we investigate the momentum distribution of the layer Berry curvature ${\cal F}_{xy}({\bf k}_{||},z)$ near the two Weyl-like points $z_{c\pm}$.
Figure 4 shows the momentum distribution of the layer Berry curvature ${\cal F}_{xy}({\bf k}_{||},z)$ in the $(k_{x},k_{y})$ plane for different layers: (a) z=10az\!=\!10a, (b) z=11az\!=\!11a, (c) z=20az\!=\!20a, and (d) z=21az\!=\!21a. Here, $a$ is the lattice constant. The two critical planes are located at $z_{c-}\!=\!10\frac{1}{3}a$ and $z_{c+}\!=\!20\frac{2}{3}a$, given by $z_{c\pm}\!=\!\frac{\lambda_{z}}{2\pi}\left[\pi\pm\arccos\left(\frac{m_{z}}{V_{z}}\right)\right]$ [see Eq. (7)], with parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$. Thus, z=10az\!=\!10a and z=11az\!=\!11a are close to the lower critical plane $z_{c-}$, while z=20az\!=\!20a and z=21az\!=\!21a are near the upper critical plane $z_{c+}$.
As shown in Fig. 4, the sign of the layer Berry curvature changes as the layer index increases from z=10az\!=\!10a [a topologically trivial region, see Fig. 1(c1)] to z=11az\!=\!11a [a topologically nontrivial region, also shown in Fig. 1(c1)]. This sign change corresponds to the band inversion occurring between these two layers, as illustrated in Fig. 1(d1). Specifically, the energy gap present at z=10az\!=\!10a closes at the lower critical plane $z\!=\!z_{c-}$ and reopens at z=11az\!=\!11a, signaling a topological phase transition.
Furthermore, the momentum distribution of the layer Berry curvature at z=11az\!=\!11a is more broadly concentrated around the center of the BZ ($k_{x}\!=\!k_{y}\!=\!0$) than that at z=10az\!=\!10a. These observations are consistent with the Chern number being 1 in the region between the two critical planes ($z_{c-}\!<\!z\!<\!z_{c+}$) and 0 outside this region ($z\!<\!z_{c-}$ and $z\!>\!z_{c+}$), as indicated in Fig. 1(c1).
A similar analysis applies to the layer Berry curvature across the upper critical plane $z\!=\!z_{c+}$.
@@ III Interlayer coupling effects
We now incorporate the effects of interlayer coupling into our analysis. The energy bands are still obtained by calculating the local density of states Chen et al. (2021); Wan et al. (2011); Yang et al. (2011), while the quantum transport properties are captured by evaluating the layer Hall conductance Fukui et al. (2005); Sun et al. (2020); Otrokov et al. (2019); Gao et al. (2021); Chen et al. (2024a); Mong et al. (2010); Essin et al. (2009); Wang et al. (2015); Varnava and Vanderbilt (2018); Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019) in the presence of interlayer coupling. Additionally, we examine how interlayer coupling influences the total 3D Hall conductivity.
@@
where $C_{j_{z}}^{{\dagger}}$ ($C_{j_{z}}$) is the electron creation (annihilation) operator on the $j_{z}$th layer, and $\eta_{z}\sigma_{z}$ represents the nearest-neighbor interlayer hopping matrix. Combining Eq. (1) with the interlayer hopping term, the full Hamiltonian incorporating interlayer coupling becomes
@@
The matrix representation of the Hamiltonian in Eq. (18) under OBCs along the $z$ direction is provided in Section SV in the Supplemental Material Sup .
@@ III.1 Local density of states
Figure 5: Local density of states (LDOS) [Eq. (19)] for different interlayer hopping strengths $\eta_{z}$. The LDOS is indicated by the red color, while the two green vertical dashed lines mark the positions of the Weyl-like points. (a) $\eta_{z}/t_{0}\!=\!0$, (b) $\eta_{z}/t_{0}\!=\!0.05$, (c) $\eta_{z}/t_{0}\!=\!0.1$, and (d) $\eta_{z}/t_{0}\!=\!0.2$. The other parameters are $k_{x}\!=\!k_{y}\!=\!0$, $k_{B}T=0.01$ eV, $V_{z}\!=\!1$ eV, $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, and $L_{z}\!=\!\lambda_{z}\!=\!31a$, with a=1a\!=\!1 nm.
@@
and $g(E,E_{n},T)$ is a Lorentzian broadening function defined as
@@
Here, $E$ is the probe energy, $k_{B}$ is the Boltzmann constant, and $T$ is the temperature.
@@
For weak interlayer coupling $\eta_{z}/t_{0}\!=\!0.05$, Fig. 5(b) reveals that the local density of states still exhibits sharp crossings at the same Weyl-like points, like the decoupled case in Fig. 5(a). For moderate coupling, $\eta_{z}/t_{0}\!=\!0.1$, Fig. 5(c) shows that the local density of states continues to display distinct crossings at the Weyl-like points, indicating that the features observed in the decoupled case persist. These results suggest that the Weyl-like points are robust against weak to moderate interlayer coupling. However, as the interlayer hopping strength $\eta_{z}$ increases further, Fig. 5(d) shows that the local density of states becomes increasingly dispersive compared with the decoupled scenario. This behavior indicates that strong interlayer coupling disrupts the localized character of the Weyl-like points, ultimately leading to their destruction.
@@ III.2 Layer Hall conductance
Figure 6: Layer Hall conductances for different interlayer hopping strengths $\eta_{z}$. The two green vertical dashed lines indicate the positions of the Weyl-like points. (a) $\eta_{z}/t_{0}\!=\!0$, (b) $\eta_{z}/t_{0}\!=\!0.05$, (c) $\eta_{z}/t_{0}\!=\!0.1$, (d) $\eta_{z}/t_{0}\!=\!0.2$, (e) $\eta_{z}/t_{0}\!=\!0.4$, and (f) $\eta_{z}/t_{0}\!=\!0.6$. All other parameters are the same as those used in Fig. 5.
@@
However, as the interlayer hopping strength $\eta_{z}$ increases further [Figs. 6(d)–6(f)], the layer Hall conductance becomes increasingly smooth and continuous in the vicinity of the Weyl-like points, deviating from the sharp features observed in the decoupled case. Physically, this behavior indicates that strong interlayer coupling destroys the Weyl-like points and significantly alters the bulk topology of the system.